1. SLOPE TAPING. A measurement is made along a line that is inclined by a vertical angle of


15°25’ as measured using a hand level and clinometer. The slope measurement is 756.52m. What


is the corresponding horizontal distance?



2. CORRECTION OF TAPE LENGTH. The correct distance between points E and F is 213.50m. If a 100m tape that is 0.025m too long is used to measure EF, what will be the measured distance?



3. CORRECTION OF TAPE LENGTH. The sides of a square lot having an area of 2.25 hectares


were measured using 100m tape that was 0.04m too short. Compute the error in the area.

1
Expert's answer
2021-10-31T00:05:11-0400

SLOPE TAPING. A measurement is made along line that is inclined by a vertical angle of 15°25’ as measured using a hand level and clinometer. The slope measurement is 756.52m. What is the corresponding horizontal distance?

s= 756.52m

α= 15°25’

d= s cosα

d= (756.52)cos(15°25’)

d = 729.30m


CORRECTION OF TAPE LENGTH. The correct distance between points E and F is 213.50m. If a 100m tape that is 0.025m too long is used to measure EF, what will be the measured distance?


NL = 100m

Corr = 0.025m

ML = 213.50m

C1= Corr(ML/NL)

= 0.025(213.50m/100m)

C1= 0.053m

CL= ML + C1

= 213.50m+ 0.053m

CL = 213.55m


CORRECTION OF TAPE LENGTH. The sides of a square lot having an area of 2.25 hectares were measured using 100m tape that was 0.04m too short. Compute the error in the area.

As= 2.25 hectares x (10000m2/1 hectare)

As= 22,500 m2

ML = √As

= √22,500m2

ML = 150m

NL = 100m

Corr = 0.04m(too short)

C1= Corr(ML/NL)

= 0.04(150m/100m)

C1= 0.06m

CL= ML-C1

= 150m -0.06m

CL = 149.94m

CAs= (149.94m)2

CAs= 22482.0036m2

Error in the area = 22,500 m2-22482.0036m2

Error in the area = 17.9964m2